# Finite Neural Networks as Mixtures of Gaussian Processes: From Provable Error Bounds to Prior Selection

### v1

@@ 5.1 Signatures: the Discretization of a Continuous Distribution
${\Delta}_{\mathcal{C}}({{z}})=\underset{{{c}}\in{\mathcal{C}}}{\mathrm{argmin}}\|{{z}}-{{c}}\|.$
The push-forward operation induced by ${\Delta}_{\mathcal{C}}$ is a mapping from ${\mathcal{P}}({\mathbb{R}}^{d})$ to ${\mathcal{D}}_{N}({\mathbb{R}}^{d})$ and defines the signature of a probability distribution. That is, as formalized in Definition 6 below, a signature induced by ${\Delta}_{\mathcal{C}}$ is an approximation of a continuous distribution with a discrete one of support with a cardinality $N$.
@@
${\mathcal{R}}_{i}=\{{{z}}\in{\mathbb{R}}^{d}:\|{{z}}-{{c}}_{i}\|\leq\|{{z}}-{{c}}_{j}\|,\forall j\in\mathbb{N}_{N},j\neq i\}.$ (6)
The intuition behind a signature is illustrated in Figure 3: the signature of distribution $p$ is a discretization of $p$ that assigns to each location ${{c}}_{i}\in{\mathcal{C}}$ a probability given by the probability mass of ${\mathcal{R}}_{i}$ according to $p$. Note that partition $\{{\mathcal{R}}_{i}\}_{i=1}^{N}$, as defined in Eqn (6), is the Voronoi partition of ${\mathbb{R}}^{d}$ w.r.t. the Euclidean distance. Hence, the signature of a probability distribution can be interpreted as a discretization of the probability distribution w.r.t. a Voronoi partition of its support. In the remainder, we refer to ${\mathcal{C}}$ as the locations of the signature, $|{\mathcal{C}}|$ as the signature size, $\{{\mathcal{R}}_{i}\}_{i=1}^{N}$ as the partition of a signature, and we call ${\Delta}$ the signature operation. In the next subsection we show how we can efficiently bound $W_{2}(p,{\Delta}_{\mathcal{C}}\#p)$ in the case where $p$ is a Gaussian mixture distribution.
@@
The approximation of a continuous distribution by a discrete distribution, also called a codebook or particle approximation, is a well-known concept in the literature (Graf and Luschgy, 2007; Ambrogioni et al., 2018; Pages and Wilbertz, 2012). The notion of the signature of a probability distribution introduced here is unique in that the discrete approximation is fully defined by the Voronoi partition of the support of the continuous distribution, thereby connecting the concept of signatures to semi-discrete optimal transport (Peyré et al., 2019).
@@ 5.2 Wasserstein Bounds for the Signature Operation
The computation of $W_{2}(p,{\Delta}_{\mathcal{C}}\#p)$ requires solving a semi-discrete optimal transport problem, where we need to find the optimal transport plan from each point ${{z}}$ to a specific location ${{c}}_{i}\in{\mathcal{C}}$ (Peyré et al., 2019). Luckily, as illustrated in the following proposition, which is a direct consequence of Theorem 1 in (Ambrogioni et al., 2018), for the case of a signature of a probability distribution, the resulting transport problem can be solved exactly.
@@
For a probability distribution $p\in{\mathcal{P}}_{2}({\mathbb{R}}^{n})$, and signature locations ${\mathcal{C}}=\{{{c}}_{i}\}_{i=1}^{N}\subset{\mathbb{R}}^{n}$, we have that
${\mathbb{W}}_{2}^{2}(p,{\Delta}_{\mathcal{C}}\#p)=\sum_{i=1}^{N}\underbrace{{{{\pi}}^{(i)}}{\mathbb{E}}_{{{z}}\sim p}[\|{{z}}-{{c}}_{i}\|^{2}\mid{{z}}\in{\mathcal{R}}_{i}]}_{\eqqcolon{\mathbb{W}}_{2\mid{\mathcal{R}}_{i}}^{2}}.$
According to Proposition 8, transporting the probability mass at each point ${{z}}$ to a discrete location ${{c}}_{i}$ based on the Voronoi partition of ${\mathbb{R}}^{d}$ guarantees that the cost $\|{{z}}-{{c}}_{i}\|$ is the smallest and leads to the smallest possible transportation cost, i.e., the optimal transportation strategy for the Wasserstein Distance. Proposition 8 is general and guarantees that for any distribution $p\in{\mathcal{P}}_{2}({\mathbb{R}}^{n})$, ${\mathbb{W}}_{2}(p,{\Delta}_{\mathcal{C}}\#p)$ only depends on ${\mathbb{W}}_{2\mid{\mathcal{R}}_{i}}$, the partial $2-$moment of $p$ w.r.t. regions ${\mathcal{R}}_{i}\in\{{\mathcal{R}}_{1},...,{\mathcal{R}}_{N}\}$. In the following proposition, we show how closed-form expressions for ${\mathbb{W}}_{2\mid{\mathcal{R}}_{i}}^{2}$ can be derived for univariate Gaussian distributions. This result is then generalized in Corollary 10 to general multivariate Gaussian distributions.
@@
${\mathbb{W}}_{2}^{2}({\Delta}_{{\mathcal{C}}}\#\mathcal{N}(0,1),\mathcal{N}(0,1))=\sum_{i=1}^{N}\nu_{i}+(\mu_{i}-c_{i})^{2},$ (7)
@@
$\mu_{i}$ $=\phi(l_{i})-\phi(u_{i}),$ (8)
$\nu_{i}$ $=1+l_{i}\phi(l_{i})-u_{i}\phi(u_{i})-[\phi(l_{i})-\phi(u_{i})]^{2}$ (9)
where $\phi$ is the pdf of a standard (univariate) Gaussian distribution, i.e., $\phi(x)=\frac{1}{\sqrt{2\pi}}\exp(-\frac{x^{2}}{2})$.
In Proposition 9, $\mu_{i}$ and $\nu_{i}$ are the mean and variance of a standard univariate Gaussian distribution restricted on $[l_{i},u_{i}]$. Note that some of the regions in the partition will be unbounded. However, as the standard Gaussian distribution exponentially decays to zero for large $x$, i.e., $\lim_{x\rightarrow\infty}\exp(x^{2})\mathcal{N}(x\mid 0,1)=0$, it follows that the bound in Eqn. (7) is finite even if some of the regions in $\{{\mathcal{R}}_{i}\}_{i=1}^{N}$ are necessarily unbounded.
A corollary of Proposition 9 is Corollary 10, where we extend Proposition 9 to general multivariate Gaussian distributions under the assumption that locations ${\mathcal{C}}$ are such that sets $\{{\mathcal{R}}_{i}\}_{i=1}^{N}$ define a grid in the transformed space induced by the basis of the covariance matrix that we denote as ${{T}}$. As illustrated in Figure 4, it is always possible to satisfy this assumption by taking ${\mathcal{C}}$ as the image of an axis-aligned grid of points under transformation ${{T}}^{-1}$, where ${{T}}^{-1}$ can be computed via an eigendecomposition (Kasim, 2020) or Cholesky decomposition (Davis et al., 2016) of the covariance matrix. [^5] Given the eigenvalues vector ${{\lambda}}$ and eigenvector matrix ${{V}}$ of a covariance matrix $\Sigma$, we can take ${{T}}^{-1}={{V}}\mathrm{diag}({{\lambda}})$. In the case of a degenerate multivariate Gaussian, where $\Sigma$ is not full rank, we can take ${\mathcal{C}}$ as the image of an axis-aligned grid of points in a space of dimension $\mathrm{rank}(\Sigma)$, under transformation ${{T}}^{-1}={({{V}}\mathrm{diag}({{\lambda}})^{\frac{1}{2}})^{{(:,1:\mathrm{rank}(\Sigma))}}}$.
@@
For $\mathcal{N}({m},\Sigma)\in\mathcal{G}({\mathbb{R}}^{n})$, let matrix ${{T}}\in{\mathbb{R}}^{n\times n}$ be such that ${{T}}=\mathrm{diag}({{\lambda}})^{-\frac{1}{2}}{{V}}^{T}$ where $\mathrm{diag}({{\lambda}})={{V}}^{T}\Sigma{{V}}$ is a diagonal matrix whose entries are the eigenvalues of $\Sigma$, and ${{V}}$ is the corresponding orthogonal (eigenvector) matrix. Further, let ${\mathcal{C}}=\{{{c}}_{i}\}_{i=1}^{N}\subset{\mathbb{R}}^{n}$ be a set of signature locations on a grid in the transformed space induced by ${{T}}$, i.e.,
$\mathrm{Post}({\mathcal{C}},{{T}})-\{{m}\}=\bar{{\mathcal{C}}}^{1}\times\bar{{\mathcal{C}}}^{2}\times\ldots\times\bar{{\mathcal{C}}}^{n}$ (10)
with $\bar{{\mathcal{C}}}^{j}=\{{{T}}^{(j,:)}({\mathcal{C}}-{m})\}\subset{\mathbb{R}}$ the set of signature locations in the transformed space for dimension $j$. Then, it holds that
${\mathbb{W}}^{2}_{2}({\Delta}_{\mathcal{C}}\#\mathcal{N}({m},\Sigma),\mathcal{N}({m},\Sigma))=\sum_{j=1}^{n}{{{\lambda}}^{(j)}}{\mathbb{W}}^{2}_{2}({\Delta}_{\bar{{\mathcal{C}}}^{j}}\#\mathcal{N}(0,1),\mathcal{N}(0,1)).$ (11)
@@
For a multivariate Gaussian $p=\mathcal{N}({m},\Sigma)$, the set ${\mathcal{C}}$ of $N>1$ signature locations that minimize ${\mathbb{W}}_{2}({\Delta}_{\mathcal{C}}\#p,p)$ is generally non-unique, and finding any such set is computationally intractable (Graf and Luschgy, 2007). However, for univariate Gaussians, an optimal signature placement strategy exists as outlined in Pagès and Printems (2003) and will be used in Subsection 7.1 to construct grid-constrained signature locations ${\mathcal{C}}$ for multivariate Gaussians.
In the case of Gaussian mixture distributions, i.e., where $p=\sum_{i=1}^{M}{\bar{\pi}^{{(i)}}}\mathcal{N}(m_{i},\Sigma_{i})$, we can simply apply a signature operation to each of the Gaussian distributions in the mixture. In particular, if we call ${\mathcal{C}}=\{{\mathcal{C}}_{1},\ldots,{\mathcal{C}}_{M}\}$ the resulting sets of locations, where ${\mathcal{C}}_{i}$ are the signature locations corresponding to $\mathcal{N}(m_{i},\Sigma_{i})$, then we have that

### v2

@@ 5.1 Signatures: the Discretization of a Continuous Distribution
${\Delta}_{\mathcal{C}}({{z}})\coloneqq\underset{{{c}}\in{\mathcal{C}}}{\mathrm{argmin}}\|{{z}}-{{c}}\|.$ (6)
The pushforward operation induced by ${\Delta}_{\mathcal{C}}$ is a mapping from ${\mathcal{P}}({\mathbb{R}}^{d})$ to ${\mathcal{D}}_{N}({\mathbb{R}}^{d})$ and defines the signature of a probability distribution. That is, as formalized in Definition 6 below, a signature induced by ${\Delta}_{\mathcal{C}}$ is an approximation of a continuous distribution with a discrete one of support with a cardinality $N$.
@@
${\mathcal{R}}_{i}\coloneqq\left\{{{z}}\in{\mathbb{R}}^{d}\colon\|{{z}}-{{c}}_{i}\|\leq\|{{z}}-{{c}}_{j}\|,\forall j\in\{1,\ldots,N\},j\neq i\right\}.$ (7)
The intuition behind a signature is illustrated in Figure 3: the signature of distribution $p$ is a discretization of $p$ that assigns to each location ${{c}}_{i}\in{\mathcal{C}}$ a probability given by the probability mass of ${\mathcal{R}}_{i}$ according to $p$. Note that partition $\{{\mathcal{R}}_{i}\}_{i=1}^{N}$, as defined in Eqn (7), is the Voronoi partition of ${\mathbb{R}}^{d}$ w.r.t. the Euclidean distance. Hence, the signature of a probability distribution can be interpreted as a discretization of the probability distribution w.r.t. a Voronoi partition of its support. In the remainder, we refer to $\mathcal{C}$} as the locations of the signature, $|{\mathcal{C}}|$ as the signature size, $\{{\mathcal{R}}_{i}\}_{i=1}^{N}$ as the partition of a signature, and we call $\Delta$} the signature operation.
@@
The signature of a probability distribution admits a closed-form expression when $\pi$ in Definition 6 is analytically tractable, i.e., the probability mass over the signature’s partition $\{{\mathcal{R}}_{i}\}_{i=1}^{N}$, is analytically tractable. For a Gaussian distribution, this is the case when the regions ${\mathcal{R}}_{i}$ are aligned with the geometry induced by the covariance matrix. In particular, for $\mathcal{N}({{\mu}},\Sigma)$, axis-aligned hyperrectangles in the eigenbasis of $\Sigma$, ellipsoids of the form $\{{{x}}\colon({{x}}-{{\mu}})^{T}\Sigma^{-1}({{x}}-{{\mu}})\leq c\}$, and unions or intersections of such sets all admit a closed-form expressions for their probability under $\mathcal{N}({{\mu}},\Sigma)$ (Genz and Bretz, 2009).
As a consequence of Remark 7, for a Gaussian mixture distribution with components with different eigenbases of their covariance, we apply the signature operation separately to each Gaussian distribution in the mixture. Specifically, for a Gaussian mixture distribution $p=\sum_{i=1}^{N}{{{\pi}}^{(i)}}p_{i}$ and a set of signature locations $\bm{{\mathcal{C}}}=\{{\mathcal{C}}_{i}\}_{i=1}^{N}$, we define the component-wise signature of $p$ as [^8] An alternative approach would be to apply the same signature locations to all components in the mixture, and approximate the resulting discrete distribution via numerical integration. However, this becomes intractable when component covariances differ significantly.
$\bm{{\Delta}}_{\bm{{\mathcal{C}}}}\#p\coloneqq\sum_{i=1}^{N}\pi^{(i)}{\Delta}_{{\mathcal{C}}_{i}}\#p_{i}.$ (8)
In the next subsection, we show how to efficiently bound ${\mathbb{W}}_{2}(p,\bm{{\Delta}}_{\bm{{\mathcal{C}}}}\#p)$.
Remark 8
The approximation of a continuous distribution by a discrete distribution, also called a particle- or quantization approximation, is a well-known concept in the literature (Graf and Luschgy, 2000; Pages and Wilbertz, 2012). The notion of the signature of a probability distribution introduced here is unique in that the discrete approximation is fully defined by the Voronoi partition of the support of the continuous distribution, thereby connecting the concept of signatures to semi-discrete optimal transport (Peyré et al., 2019).
@@ 5.2 Wasserstein Bounds for the Signature Operation
The computation of $W_{2}(p,{\Delta}_{\mathcal{C}}\#p)$ requires solving a semi-discrete optimal transport problem, where we need to find the optimal transport plan from each point z{{z}} to a specific location ${{c}}_{i}\in{\mathcal{C}}$ (Peyré et al., 2019). Luckily, as illustrated in the following proposition, which is a direct consequence of Lemma 3.1 in (Canas and Rosasco, 2012), for the case of a signature of a probability distribution, the resulting transport problem can be solved exactly.
@@
Let $\rho\geq 1$, then for a probability distribution $p\in{\mathcal{P}}_{\rho}({\mathbb{R}}^{n})$, and signature locations ${\mathcal{C}}=\{{{c}}_{i}\}_{i=1}^{N}\subset{\mathbb{R}}^{n}$, we have that
${\mathbb{W}}{{}_{\rho}^{\rho}}(p,{\Delta}_{\mathcal{C}}\#p)=\sum_{i=1}^{N}{{{\pi}}^{(i)}}{\mathbb{E}}_{{{z}}\sim p}[\|{{z}}-{{c}}_{i}\|^{\rho}\mid{{z}}\in{\mathcal{R}}_{i}].$ (9)
According to Proposition 9, transporting the probability mass at each point z{{z}} to a discrete location ${{c}}_{i}$ based on the Voronoi partition of ${\mathbb{R}}^{d}$ guarantees that the cost $\|{{z}}-{{c}}_{i}\|$ is the smallest and leads to the smallest possible transportation cost, i.e., the optimal transportation strategy for the Wasserstein Distance. Proposition 9 is general and guarantees that for any distribution $p\in{\mathcal{P}}_{2}({\mathbb{R}}^{n})$, ${\mathbb{W}}_{2}(p,{\Delta}_{\mathcal{C}}\#p)$ only depends on the probability mass and the conditional $2$-moment of $p$ w.r.t. regions ${\mathcal{R}}_{i}\in\{{\mathcal{R}}_{1},...,{\mathcal{R}}_{N}\}$. In the following proposition, we show how closed-form expressions for ${\mathbb{W}}_{2}^{2}(p,{\Delta}_{\mathcal{C}}\#p)$ can be derived for univariate Gaussian distributions. This result is then generalized in Corollary 11 to general multivariate Gaussian distributions.
@@
${\mathbb{W}}_{2}^{2}({\Delta}_{{\mathcal{C}}}\#\mathcal{N}(0,1),\mathcal{N}(0,1))=\sum_{i=1}^{N}{{{\pi}}^{(i)}}\left[\nu_{i}+(\mu_{i}-c_{i})^{2}\right]$ (10)
@@
${{{\pi}}^{(i)}}=\Phi(u_{i})-\Phi(l_{i}),\quad\mu_{i}=\frac{1}{{{{\pi}}^{(i)}}}[\phi(l_{i})-\phi(u_{i})],\quad\nu_{i}=1+\frac{1}{{{{\pi}}^{(i)}}}[l_{i}\phi(l_{i})-u_{i}\phi(u_{i})]-\mu_{i}^{2}$ (11)
where $\phi$ and $\Phi$ are the pdf and cdf of a standard (univariate) Gaussian distribution, respectively, i.e., $\phi(x)=\tfrac{1}{\sqrt{2\pi}}\exp\left(-\tfrac{x^{2}}{2}\right)$ and $\Phi(x)=\tfrac{1}{2}\left[1+\mathrm{erf}\left(\tfrac{x}{\sqrt{2}}\right)\right]$.
In Proposition 10, $\mu_{i}$ and $\nu_{i}$ are the mean and variance of a standard univariate Gaussian distribution restricted on $l_{i},u_{i}]$. Note that some of the regions in the partition will be unbounded. However, as the standard Gaussian distribution exponentially decays to zero for large $x$, i.e., $\lim_{x\rightarrow\infty}\exp(x^{2})\mathcal{N}(x\mid 0,1)=0$, it follows that the bound in Eqn (10) is finite even if some of the regions in $\{{\mathcal{R}}_{i}\}_{i=1}^{N}$ are necessarily unbounded.
Figure 4: Construction of the signature of a 2D Gaussian distribution (black dots) using the signature of a standard Gaussian distribution (blue dots) as a template. In the space induced by T{{T}}, that is, the basis of the covariance matrix (orange arrows) of the Gaussian distribution, all dimensions of the Gaussian distribution are independent. Hence, we can take the cross-product of the signatures of each univariate Gaussian marginal in dimension $j$ of the transformed space with signature locations ${\mathcal{C}}^{j}$ (green stripes) as a template. The signature in the original space is then obtained by taking the post image of the template under ${{T}}^{-1}$. The blue and black lines represent the edges of the Voronoi partition associated with the signature locations in the transformed and original space, respectively.
A corollary of Proposition 10 is Corollary 11, where we extend Proposition 10 to general multivariate Gaussian distributions under the assumption that locations $\mathcal{C}$} are such that sets $\{{\mathcal{R}}_{i}\}_{i=1}^{N}$ define a grid in the transformed space induced by the basis of the covariance matrix that we denote as T{{T}}. As illustrated in Figure 4, it is always possible to satisfy this assumption by taking $\mathcal{C}$} as the image of an axis-aligned grid of points under transformation ${{T}}^{-1}$, where ${{T}}^{-1}$ can be computed via an eigendecomposition (Kasim, 2020) or Cholesky decomposition (Davis et al., 2016) of the covariance matrix. [^9] Given the eigenvalues vector $\lambda$}} and eigenvector matrix V{{V}} of a covariance matrix $\Sigma$, we can take ${{T}}^{-1}={{V}}\mathrm{diag}({{\lambda}})$. In the case of a degenerate multivariate Gaussian, where $\Sigma$ is not full rank, we can take $\mathcal{C}$} as the image of an axis-aligned grid of points in a space of dimension $\mathrm{rank}(\Sigma)$, under transformation ${{T}}^{-1}={({{V}}\mathrm{diag}({{\lambda}})^{\frac{1}{2}})^{{(:,1:\mathrm{rank}(\Sigma))}}}$.
@@
For $\mathcal{N}({m},\Sigma)\in\mathcal{G}({\mathbb{R}}^{n})$, let matrix ${{T}}\in{\mathbb{R}}^{n\times n}$ be such that ${{T}}=\mathrm{diag}({{\lambda}})^{-\frac{1}{2}}{{V}}^{T}$ where $\mathrm{diag}({{\lambda}})={{V}}^{T}\Sigma{{V}}$ is a diagonal matrix whose entries are the eigenvalues of $\Sigma$, and V{{V}} is the corresponding orthogonal (eigenvector) matrix. Further, let ${\mathcal{C}}=\{{{c}}_{i}\}_{i=1}^{N}\subset{\mathbb{R}}^{n}$ be a set of signature locations on a grid in the transformed space induced by T{{T}}, i.e.,
${\mathrm{Post}({\mathcal{C}}-\{{m}\},{{T}})}={\mathcal{C}}^{1}\times{\mathcal{C}}^{2}\times\ldots\times{\mathcal{C}}^{n}$ (12)
with ${\mathcal{C}}^{j}=\{c_{i}\}_{i=1}^{N_{j}}$ the set of signature locations in the transformed space for dimension $j$. Then,
${\mathbb{W}}^{2}_{2}({\Delta}_{\mathcal{C}}\#\mathcal{N}({m},\Sigma),\mathcal{N}({m},\Sigma))=\sum_{j=1}^{n}{{{\lambda}}^{(j)}}{\mathbb{W}}^{2}_{2}\left({\Delta}_{{\mathcal{C}}^{j}}\#\mathcal{N}(0,1),\mathcal{N}(0,1)\right).$ (13)
In the case of Gaussian mixture distributions, i.e., where $p=\sum_{i=1}^{M}{\tilde{\pi}^{{(i)}}}\mathcal{N}(m_{i},\Sigma_{i})$, and the component-wise signature operation defined in Eqn (8), i.e., $\bm{{\Delta}}_{\bm{{\mathcal{C}}}}$ with $\bm{{\mathcal{C}}}=\{{\mathcal{C}}_{i=1}^{N}\}$, we have that
${\mathbb{W}}_{2}^{2}\left(p,\bm{{\Delta}}_{\bm{{\mathcal{C}}}}\#p\right)\leq\sum_{i=1}^{M}{\tilde{\pi}^{{(i)}}}{\mathbb{W}}_{2}^{2}\left(p_{i},{\Delta}_{{\mathcal{C}}_{i}}\#p_{i}\right)\eqqcolon\hat{\mathbb{W}}_{2,\bm{{\Delta}}_{\bm{{\mathcal{C}}}}}^{2}(p).$ (14)
That is, for a mixture of Gaussian distributions, the $2$-Wasserstein distance between the mixture and its component-wise signature is bounded by the weighted sum of the 2-Wasserstein distances between each Gaussian component and its signature. [^10] This trivial result is a special case of Lemma 31 in the Appendix on the Wasserstein distance between mixtures. Note that, under the assumption that the signature locations of each component satisfy Eqn (12), this upper bound admits a closed form as given by Corollary 11.
@@
For a multivariate Gaussian $p=\mathcal{N}({m},\Sigma)$, the set $\mathcal{C}$} of $N>1$ signature locations that minimize ${\mathbb{W}}_{2}({\Delta}_{\mathcal{C}}\#p,p)$ is generally non-unique, and finding any such set is computationally intractable (Graf and Luschgy, 2000). However, for univariate Gaussians, an optimal signature placement strategy exists as outlined in Pagès and Printems (2003) and will be used in Subsection 7.1 to construct grid-constrained signature locations $\mathcal{C}$} for multivariate Gaussians.
5.3 Convergence Rates for Signature Approximations of Gaussian Mixtures
In the previous section, we showed how we can obtain a closed form expression for ${\mathbb{W}}_{2}(p,{\Delta}_{\mathcal{C}}\#p)$ when $p$ is a multivariate Gaussian distribution, and that this expression can be used to upper bound the $2$-Wasserstein distance error from the component-wise signature $\bm{{\Delta}}_{\bm{{\mathcal{C}}}}\#p$ when $p$ is a mixture of Gaussian distributions. Here, we show that this error converges uniformly to zero as the number of signature locations increases. We start with the case where $p$ is a univariate Gaussian distribution, and then extend to mixtures of Gaussian distributions. To our knowledge, the resulting uniform, non‑asymptotic convergence rates are novel. Indeed, existing works on quantization for Gaussian measures (e.g., Graf and Luschgy (2000); Pagès and Printems (2003)) typically provide asymptotic rates and existence results for optimal $N$-point quantizers, rather than the explicit uniform‑in‑$N$ error bounds derived here.
Proposition 13
For $p=\mathcal{N}(0,1)$, consider $N\in\mathbb{N}$ signature locations on a uniform grid centered at zero with spacing $(2\sqrt{\log N})/N$ between consecutive points, [^11] We choose the spacing $(2\sqrt{\log N})/N$ as it balances the discretization error in the central interval and the Gaussian tail mass. Other spacings trade off these contributions differently: for example, a spacing of $(\log N)/N$ would reduce the remainder term $r(N)$ but increase the leading term to $(\log^{2}N)/N^{2}$. that is
${\mathcal{C}}=\left\{\frac{2\sqrt{\log N}}{N}\left(2i-N-1\right)\colon i\in\{1,\ldots,N\}\right\}.$ (15)
Then, for all $N\in\mathbb{N}$,
${\mathbb{W}}_{2}^{2}({\Delta}_{\mathcal{C}}\#\mathcal{N}(0,1),\mathcal{N}(0,1))\leq\frac{4\log N}{N^{2}}+r(N),$ (16)
where the remainder term
$r(N)=\begin{cases}1,&N=1,\\ \frac{4}{N^{2}\sqrt{2\pi}}\left(\sqrt{\log N}+\frac{1}{4\sqrt{\log N}}\right),&N\geq 2.\end{cases}$ (17)
The proof of Proposition 13 (given in the Appendix) follows by applying Proposition 10 to write the ${\mathbb{W}}_{2}$ error as a sum over the Voronoi partition induced by $\mathcal{C}$}, then bound each interval’s contribution by its squared radius and control the unbounded intervals using Gaussian tail bounds (Mills’ ratio).
Remark 14
For $N\geq 2$, the remainder term of the bound of Proposition 13 (Eqn (17)) is
$r(N)=\tfrac{4\sqrt{\log N}}{N^{2}\sqrt{2\pi}}+O\!\left(\tfrac{1}{N^{2}\sqrt{\log N}}\right).$
Hence $r(N)=O\!\left(\tfrac{\sqrt{\log N}}{N^{2}}\right)$ and $r(N)=o\!\left(\tfrac{\log N}{N^{2}}\right)$, so the bound in Eqn (16) is dominated by its first term $\tfrac{4\log N}{N^{2}}$, and consequently
${\mathbb{W}}_{2}^{2}({\Delta}_{\mathcal{C}}\#\mathcal{N}(0,1),\mathcal{N}(0,1))\leq\frac{4\log N}{N^{2}}\bigl{(}1+o(1)\bigr{)}.$
That is, for a uniform grid, the squared $W_{2}$ error decays on the order of $(\log N)/N^{2}$. In one dimension, Theorem 6.2 of Graf and Luschgy (2000) implies that the minimal achievable squared ${\mathbb{W}}_{2}$ error of any $N$-location signature approximating a univariate Gaussian decays as $1/N^{2}$ up to constants. Thus the uniform grid in Proposition 13 attains the optimal polynomial rate $1/N^{2}$ up to a single logarithmic factor. The $\log N$ term arises from using a uniform construction that provides an explicit bound for all $N$, with a remainder of strictly lower order than the leading term. In practice, adaptive (non-uniform) placement (Remark 12) can remove this logarithmic factor and improve the constants in the error bound, as illustrated in Figure 5.
(a) Formal upper bounds on ${\mathbb{W}}_{2}$ for uniform and adaptive grid-constrained signatures.
(b) Empirical ${\mathbb{W}}_{2}$ comparisons of grid‑constrained and approximate optimal signatures.
Figure 5: The 2-Wasserstein distance between a 2D Gaussian mixture distribution with $M$ components and signature approximations with NN locations. Solid lines correspond to signatures with non-uniform grid‑constrained locations obtained via Algorithm 2, while dashed lines correspond to signatures with uniform grid-constrained locations as in Corollary 15. Dash-dotted lines correspond to signatures with locations given by the cluster centers obtained via NN-Means on $10^{5}$ Monte Carlo samples from the GMM, which provide an approximation to the optimal signature locations (Pagès and Printems, 2003). For grid‑constrained signatures, formal bounds are computed using Algorithm 2, while empirical estimates are computed from $10^{4}$ samples from each distribution by computing the 2‑Wasserstein distance between the resulting empirical distributions. The dashed reference curves show the convergence rate from Corollary 15, which corresponds to the uniform grid. For comparability, each mixture has an expected squared $\ell_{2}$ norm of one, so the reported distance equals the relative 2‑Wasserstein distance defined in Eqn (34).
Analogously to how Proposition 10 extends to mixtures of Gaussian distributions, Proposition 13 extends to mixtures of multivariate Gaussian distributions, as shown in the following corollary.
Corollary 15 (of Proposition 13)
Let $p=\sum_{i=1}^{M}{\tilde{\pi}^{{(i)}}}\mathcal{N}(m_{i},\Sigma_{i})$. For each component $i$ of $p$, define a set of signature locations ${\mathcal{C}}_{i}\in{\mathbb{R}}^{n}$ as in Eqn (12), using uniform one-dimensional grid of points ${\mathcal{C}}_{i}^{j}$ of size $N_{i,j}$ as in Eqn (15) for each dimension $j$.